# Cross-Subject Data Splitting for Brain-to-Text Decoding

### v3

@@ Abstract
Recent major milestones have successfully decoded non-invasive brain signals (e.g. functional Magnetic Resonance Imaging (fMRI) and electroencephalogram (EEG)) into natural language. Despite the progress in model design, how to split the datasets for training, validating, and testing still remains a matter of debate. Most of the prior researches applied subject-specific data splitting, where the decoding model is trained and evaluated per subject. Such splitting method poses challenges to the utilization efficiency of dataset as well as the generalization of models. In this study, we propose a cross-subject data splitting criterion for brain-to-text decoding on various types of cognitive dataset (fMRI, EEG), aiming to maximize dataset utilization and improve model generalization. We undertake a comprehensive analysis on existing cross-subject data splitting strategies and prove that all these methods suffer from data leakage, namely the leakage of test data to training set, which significantly leads to overfitting and overestimation of decoding models. The proposed cross-subject splitting method successfully addresses the data leakage problem and we re-evaluate some SOTA brain-to-text decoding models as baselines for further research.
@@ 1 Introduction
Brain-to-text decoding aims to recover natural language from brain signals stimulated by corresponding speech. Recent studies (Makin et al., 2020; Wang and Ji, 2022; Xi et al., 2023; Tang et al., 2023; Duan et al., 2024) have successfully decoded non-invasive brain signals (e.g. fMRI, EEG) to text by applying deep computational neural networks. However, no consensus has reached on how to split the cognitive dataset for training, validating, and testing. Most of the prior work (Ye et al., 2023; Tang et al., 2023) performed subject-specific data splitting for training and evaluating decoding models. Under this splitting rule, data for training, validating, and testing all comes from one specific subject from the cognitive dataset. For example, Tang et al. (2023) picked three subjects out of seven and conducted model training and evaluation on the three subjects respectively. This subject-specific splitting method causes two main problems. First, it only utilizes a tiny part of the whole dataset. Since the collection of brain signals is costly and time-consuming, such splitting method results in significant waste of data resources. Second, it leads to the poor generalization of decoding models. As every subject’s brain has unique functional and anatomical structures, subject-specific models may exhibit considerable variability across individuals and fail to generalize to other subjects. Moreover, decoding models trained from scratch on limited data are prone to facing the overfitting problem.
Figure 1: Illustration of naturalistic language comprehension dataset for brain-to-text decoding. Path with the same color indicates one sample for training/validating/testing.
Some studies (Wang and Ji, 2022; Xi et al., 2023) began to shed light on cross-subject data splitting, which views all the subjects’ data as a whole and performs splitting according to a given ratio (e.g. 8:1:1 for the training set, validation set, and test set). Cross-subject data splitting effectively compensates for the shortcomings of subject-specific splitting and has been widely applied in brain-to-image decoding (Wang et al., 2024; Liu et al., 2024). However, unlike datasets for brain-to-image decoding, where subjects are guided to see different and unrepeated pictures, different subjects will hear the same story in the naturalistic language comprehension dataset for brain-to-text decoding, which challenges cross-subject data splitting. As shown in Figure 1, such dataset is usually formatted in subject-task-text-signal ($S$-$M$-$T$-$F$) pair, indicating the brain signal $F$ of subject $S$ stimulated by hearing text $T$ from task $M$. Current cross-subject data splitting methods (Wang and Ji, 2022; Xi et al., 2023) can be summarized as five categories: (1) split by subjects $S$, (2) split by tasks $M$, (3) split by randomly picking signal frames $F$, (4) split by randomly picking signal frames under certain task $F$-$M$, (5) split by randomly picking consecutive signal frames under certain task $F$-$M$. However, based on our observations, all these splitting methods suffer from data leakage problem, namely part of the test data is mixed into the training set, which leads to overfitting in model training and overestimation in model evaluation.
@@
Specifically, modern brain-to-text decoding models follow an encoder-decoder manner. We pick two representative models: EEG2Text (Wang and Ji, 2022) and UniCoRN (Xi et al., 2023) for investigating the damage of data leakage. The former is an end-to-end encoder-decoder framework, while the latter first pre-trains the encoder and then applies it in the decoder training. Experiments support that data leakage affects model training on both the encoder side and decoder side. For the encoder component, if subjects’ brain signals in the test set are mixed into the training set, the encoder will become overfitting and fail to well represent unseen subjects’ brain signals. As to the decoder, the situation gets worse if text stimuli are leaked. Since the decoder follows an auto-regressive manner and generates token by token, data leakage will cause the decoder to memorize seen paragraphs during the teacher-forcing training stage, which leads to poor generalization to unseen text.
@@
To avoid data leakage and fairly evaluate the performance of brain-to-text decoding models, we propose a cross-subject data splitting criterion. We focus on fMRI and EEG signals in this study, although the proposed criterion could be applied to any datasets satisfying the prescribed format. In the proposed method, the dataset is split according to subject-text ($S$-$T$) pairs with the following rules: (1) Brain signals collected from specific subject in validation set and test set will not appear in training set, which means the trained encoder cannot get access to any brain information belonging to subjects in test set. (2) Text stimuli in validation set and test set will not appear in training set. The decoder learns to reconstruct language with brain signals instead of memorizing seen text.
@@
  • To the best of our knowledge, we propose the first cross-subject data splitting criterion for brain-to-text decoding.
  • We comprehensively analyze current cross-subject data splitting methods and find that all existing methods suffer from data leakage problem, which severely affects the training and evaluation of decoding models.
@@
  • Some SOTA brain-to-text decoding models are re-evaluated under the proposed cross-subject data splitting method as baselines for further research.

### v4

# Rethinking Cross-Subject Data Splitting for Brain-to-Text Decoding

@@ Abstract
Recent major milestones have successfully reconstructed natural language from non-invasive brain signals (e.g. functional Magnetic Resonance Imaging (fMRI) and Electroencephalogram (EEG)) across subjects. However, we find current dataset splitting strategies for cross-subject brain-to-text decoding are wrong. Specifically, we first demonstrate that all current splitting methods suffer from data leakage problem, which refers to the leakage of validation and test data into training set, resulting in significant overfitting and overestimation of decoding models. In this study, we develop a right cross-subject data splitting criterion without data leakage for decoding fMRI and EEG signal to text. Some SOTA brain-to-text decoding models are re-evaluated correctly with the proposed criterion for further research.
11footnotetext: Work done during internship in JD.com.22footnotetext: Corresponding author.
@@ 1 Introduction
Brain-to-text decoding aims to recover natural language from brain signals stimulated by corresponding speech. Recent studies (Makin et al., 2020; Wang and Ji, 2022; Xi et al., 2023; Tang et al., 2023; Duan et al., 2024) have successfully decoded non-invasive brain signals (e.g. fMRI, EEG) to text by applying deep neural networks. Most of these works perform within-subject data splitting for training and evaluating decoding models. This subject-specific splitting method causes two main problems. First, it only uses data from one subject of the whole dataset for training and testing. Since brain signal collection is costly and time-consuming, such splitting method results in a significant waste of data resources. Second, it leads to poor model generalization. As every brain has unique functional and anatomical structures, subject-specific models may exhibit considerable variability across individuals and fail to generalize to other subjects (Liu et al., 2024). Moreover, decoding models trained from scratch on limited data are prone to facing the overfitting problem.
Human brain responds similarly to the same stimuli, despite the individual discrepancy (Hasson et al., 2004; Pereira et al., 2018). Therefore, some studies (Wang and Ji, 2022; Xi et al., 2023; Duan et al., 2024) begin to shed light on cross-subject brain-to-text decoding, which performs data splitting based on all the subjects, trains and evaluates decoding model once. Cross-subject data splitting effectively compensates for the shortcomings of subject-specific splitting, and has been widely applied in brain-to-image decoding (Wang et al., 2024; Liu et al., 2024). However, unlike datasets for brain-to-image decoding (Allen et al., 2022; Chang et al., 2019) where subjects are guided to see different and unrepeated pictures, different subjects will be stimulated by the same story in common naturalistic language comprehension dataset, which challenges cross-subject data splitting.
@@
Based on our observations, current cross-subject data splitting methods for brain-to-text decoding are wrong because data for validation and test leaks into the training set, rendering the evaluation of the decoding process meaningless. Specifically, we find two types of data leakage: brain signal leakage and text stimuli leakage. Brain signal leakage refers to test subject’s brain signal appears in training set. Text stimuli leakage refers to text in test set appears in the training set. Modern brain-to-text decoding models follow an encoder-decoder manner. We pick two representative models: EEG2Text (Wang and Ji, 2022) and UniCoRN (Xi et al., 2023) to reveal data leakage and its damage. Experiments support that data leakage affects model training on both encoder and decoder side. For the encoder, it will become overfitting and fail to well represent brain signals if brain signal leakage exists. For the decoder, the situation gets worse if text stimuli leakage happens. Any data leakage would cause the auto-regressive decoder to memorize previously seen paragraphs during training stage, resulting in poor generalization to unseen text.
@@
To avoid data leakage and fairly evaluate the performance of cross-subject brain-to-text decoding models, we propose a right data splitting method. We focus on fMRI and EEG signals in this study, although the proposed criterion could be applied to any datasets satisfying the prescribed format. In the proposed method, we follow two basic rules: (1) Brain signals collected from specific subject in validation set and test set will not appear in training set, which means the trained encoder cannot get access to any brain information belonging to subjects in test set. (2) Text stimuli in validation set and test set will not appear in training set. The decoder learns to reconstruct language with brain signals instead of memorizing seen text.
@@
  • To the best of our knowledge, we are the first to identify the issue of data leakage in current cross-subject data splitting methods for brain-to-text decoding.
  • We define the splitting criterion for cross-subject brain-to-text decoding, and propose a right dataset splitting method.
@@
  • Some SOTA brain-to-text decoding models are re-evaluated using the proposed cross-subject data splitting method to ensure a fair assessment of their performance.
2 Preliminary
2.1 Dataset Description
A naturalistic language comprehension dataset $\mathcal{D}$ contains brain signals of $N$ subjects when they passively listen to $K$ spoken stories. Suppose that not all subjects are stimulated by all stories, and different subjects may hear the same story.
Formally, $S_{1},S_{2},\ldots,S_{N}$ denotes to the $N$ subjects and $M_{1},M_{2},\ldots,M_{K}$ denotes to the $K$ stories in dataset. The $k$-th story $M_{k}$ consists of $l_{k}$ text segments $T_{k1},T_{k2},\ldots,T_{kl_{k}}$. If the $i$-th subject $S_{i}$ hears the $j$-th text segment $T_{kj}$, then his brain signal is denoted as $F_{ijk}$.
Figure 1: Illustration of how to build graph to describe dataset step by step.
2.2 Use Graph to Describe Dataset
We use multigraph and $k$-partite graph (detailed in D.1 & D.2) to describe the intricate structure of naturalistic language comprehension dataset.
Definition 2.1.
A naturalistic language comprehension dataset $\mathcal{D}$ can be represented via a directed $4$-partite multigraph $\mathcal{G}_{\mathcal{D}}$.
How to build the directed $4$-partite multigraph $\mathcal{G}_{\mathcal{D}}$ step by step is shown in Figure 1. Graph 1 is a $2$-partite graph indicating subject $S_{i}$ listening to story $M_{k}$. Subject $S_{i}$ and story $M_{k}$ are viewed as vertices, and edges connecting them indicate certain type of relationship (e.g. $S_{i}$ “listen to” $M_{k}$ in this case). Graph 2 illustrates that story $M_{k}$ consists of text segments $T_{kj}$. Graph 3 shows the brain signals $F_{ijk}$ of subject $S_{i}$ stimulated by text segment $T_{kj}$. Graph 4 is an example of combining the three $2$-partite graphs Graph 1-3: $F_{122},F_{222},F_{322},F_{422}$ are brain signals of $S_{1},S_{2},S_{3},S_{4}$ stimulated by text segment $T_{22}$ from story $M_{2}$. In this example, four edges between $M_{2}$ and $T_{22}$ correspond to the different responses of four subjects to the same text segment. There are three edges between $S_{2}$ and $M_{2}$ because $M_{2}$ contains three text segments. Edges of the same color indicate one sample in dataset. Graph 5 shows the complete directed $4$-partite multigraph $\mathcal{G}_{\mathcal{D}}$ for representing whole dataset. Every sample in dataset can be represented through ordered subject-story-text-brain ($S_{i},M_{k},T_{kj},F_{ijk}$) pair. We introduce the formal notation of $\mathcal{G}_{\mathcal{D}}$: